# Designing ReLU Generative Networks to Enumerate Trees with a Given Tree Edit Distance

### v1

@@ Abstract
The generation of trees with a specified tree edit distance has significant applications across various fields, including computational biology, structured data analysis, and image processing. Recently, generative networks have been increasingly employed to synthesize new data that closely resembles the original datasets. However, the appropriate size and depth of generative networks required to generate data with a specified tree edit distance remain unclear. In this paper, we theoretically establish the existence and construction of generative networks capable of producing trees similar to a given tree with respect to the tree edit distance. Specifically, for a given rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from an alphabet $\Sigma$, and a non-negative integer $d$, we prove that all rooted, ordered, and vertex-labeled trees over $\Sigma$ with tree edit distance at most $d$ from $T$ can be generated using a ReLU-based generative network with size $\mathcal{O}(n^{3})$ and constant depth. The proposed networks were implemented and evaluated for generating trees with up to 21 nodes. Due to their deterministic architecture, the networks successfully generated all valid trees within the specified tree edit distance. In contrast, state-of-the-art graph generative models GraphRNN and GraphGDP, which rely on non-deterministic mechanisms, produced significantly fewer valid trees, achieving validation rates of only up to 35% and 48%, respectively. These findings provide a theoretical foundation towards construction of compact generative models and open new directions for exact and valid tree-structured data generation. An implementation of the proposed networks is available at https://github.com/MGANN-KU/TreeGen_ReLUNetworks.
@@ 1 Introduction
As a step towards addressing these issues we study the exact and deterministic generation of all rooted, ordered, and vertex-labeled trees similar to a given tree using neural networks and prove the existence of ReLU-based generative networks for tree edit distance. Given a rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from a symbol set $\Sigma$, we theoretically establish the existence of ReLU-activated generative networks capable of producing all rooted trees over $\Sigma$ within tree edit distance at most $d$ from $T$. The key idea of our approach is to first construct a directed, rooted, ordered, and edge-labeled tree based on $T$. We then reduce the tree edit distance problem to a string edit distance problem by representing the tree as an Euler string [48], obtained through a depth-first search (DFS) traversal. The proposed networks are applied on trees with up to 21 nodes to generate similar trees, and are compared with the state-of-the-art graph generative models GraphRNN by You et al. [53] and GraphGDP by Huang et al. [57]. An implementation of the proposed networks is available at https://github.com/MGANN-KU/TreeGen_ReLUNetworks.
The paper is organized as follows: Preliminaries are discussed in Section 2. ReLU generative networks that can identify the indices and labels of the directed edges in the Euler string are discussed in Section 3. Existence of ReLU networks to generate all trees with tree edit distance at most $d$ due to substitution operations is discussed in Section 4. Existence of ReLU networks to generate all trees with edit distance at most (resp., exactly) $d$ due to deletion (resp., insertion) operations is discussed in Section 5. Generation of all trees with tree edit distance at most $d$ due to simultaneous application of deletion, substitution and insertion operations by using a ReLU network is discussed in Section 6. Computational experiments are discussed in Section 7. A conclusion and future directions are given in Section 8. Proofs of some theorems, examples and explanations of the program codes with sample instances are given in Appendix 9.
@@ 2 Preliminaries
Observe that a tree can be completely determined by its Euler string, i.e., $E(T)$ is a canonical representation of $T$ when the labels of roots of the underlying trees is fixed. Therefore any tree edit operations (substitution, deletion, and insertion) on a non-root vertex $u$ of a given tree can be viewed as edit operations on the Euler string of the tree on the entries that correspond to the inward edge $(u,v)$ and the outward edge $(v,u)$ with some refinements to obtain the desired tree. Furthermore, a vertex in $T$ can be specified by its DFS index. Therefore in the rest of the paper, we will use a random sequence $x_{1},x_{2},\ldots,x_{d}$, $d\geq 1$, with integers $x_{j}\in[0,n]$, unless stated otherwise, to specify the DFS indices of the vertices under consideration in a tree with $n+1$ vertices. We ignore $x_{j}=0$ and repeated entries in the case of substitution and deletion operations. An example random sequence is given in Fig. 2(a). In the rest of the discussion, we call a vertex-labeled or edge-labeled, rooted and ordered tree simply a tree. A list of symbols, variables, and their descriptions used throughout the discussion is provided in Table 3.
@@ 3 Identification of Edge Labels by ReLU Network
We focus on designing generative networks with ReLU as an activation function to generate all trees that are similar to a given tree. More precisely, we are interested in the following problem:
@@
Output: Construct generative networks with ReLU as an activation function that can generate all rooted, ordered, and vertex-labeled trees over $\Sigma$ with tree edit distance at most $d$ from $T$.
@@
Proposition 1 gives a necessary and sufficient condition for the $i$-th entry to be the outward edge of an inward edge at $\ell$-th position in an Euler string. The condition essentially depends on the number of inward edges and outward edges between the two given positions $i$ and $\ell$. Before going into the details, for any two positions i{i} and $\ell$}, with $1\leq\ell<i\leq 2n$, we denote by $\rm{in}_{\ell i}$ (resp., $\rm{out}_{\ell i}$), the number of inward edges (resp., outward edges) $t_{k}$, $\ell<k<i$. Consider the tree $T$ given in Fig. 2(a) and its Euler string $E(T)$ given in Fig. 2(c), for $i=7$, $\ell=2$, $\rm{in}_{2,7}=2$ as there are two inward edges $t_{4},t_{5}$ and $\rm{out}_{2,7}=2$ as there are two outward edges $t_{3},t_{6}$ (see Fig. 4(a)). So in this case $\rm{in}_{2,7}=\rm{out}_{2,7}$. Similarly, for $i=9,\ell=3$, $\rm{in}_{3,9}=3>2=\rm{out}_{3,9}$ (see Fig. 4(b)) and for $i=9,\ell=5$, $\rm{in}_{5,9}=1<2=\rm{out}_{5,9}$ (see Fig. 4(c)).
Figure 4: An illustration of the number of inward edges (green) (resp., outward edges (blue)) for the edges $t_{i},i=7,9$ and $t_{\ell},\ell=2,3,5$ which are depicted in gray boxes.
@@
Consider the tree $T$ given in Fig. 2(a) and its Euler string $E(T)$ shown in Fig. 2(c). We want to determine if the edges $t_{i}$, $i=4,7,9$ in $E(T)$ are outward edges of some inward edges by using Proposition 1. We see that Proposition 1(ii) does not hold for $i=4$ and any $\ell\in[1,3]$, therefore $t_{4}$ is not an outward edge of any inward edge, which is consistent with the fact. For $i=7$, Proposition 1(ii) and (iii) are satisfied for both $\ell=2,4$. For $\ell=2$ (resp., $\ell=4$), $\rm{in}_{2,7}=2=\rm{out}_{2,7}$ (resp., $\rm{in}_{4,7}=1=\rm{out}_{4,7}$) (see Fig. 4(a)). That is $t_{2}$ and $t_{4}$ are both candidate inward edges of $t_{7}$. By using Proposition 1(iv), $t_{7}$ is the outward edge of $t_{4}$. For $i=9$, we see that $t_{5}$ and $t_{8}$ are the only edges that satisfy Proposition 1(i)-(ii), but $t_{5}$ does not satisfy the Proposition 1(iii). Therefore $t_{8}$ is the inward edge of $t_{9}$ implying that $t_{9}$ is an outward edge.
@@ 4 TSd-generative ReLU
Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)=t_{1},t_{2},\ldots,t_{2n}$, and a non-negative integer $d$. We define the TSd-generative ReLU to be a ReLU neural network with $2d$ input nodes $x=x_{1},x_{2},\ldots,x_{2d}$ over $\{0,\ldots,n\}$, and $2n$ output nodes $u=u_{1},u_{2},\ldots,u_{2n}$ over $\Sigma$ such that all Euler strings $u$ of trees with the tree edit distance at most $2d$ from $E(T)$ can be obtained by the substitution of appropriate $x_{j+d}$ and $x_{j+d}+m$ at the inward edge and outward edge of $x_{j}\neq 0$, respectively. In this context $x_{1},\ldots,x_{d}$ represents the inward edges for the substitution operations, while $x_{d+1},\ldots,x_{2d}$ denotes the values to be substituted during these operations. An illustration of such a network is given in Fig. 6, where $m=5$, $d=3$, $x=1,3,0,5,1,2$, means that the labels of the inward edges (resp., outward edges) of $x_{1}=1,x_{2}=3$ are substituted by $x_{4}=5,x_{5}=1$ (resp., $10,6$). The entry $x_{3}=0$ corresponds to the root, and is ignored.
@@ 5 TDd, TId-generative ReLU
Let $T$ be a tree with $n+1$ nodes, labels from $\Sigma=\{1,2,\ldots,m\}$ and Euler string $E(T)=t_{1},t_{2},\ldots,t_{2n}$, and a non-negative integer $d$. We define the TDd-generative ReLU to be a ReLU neural network that can generate all Euler strings over $\Sigma$ with the tree edit distance at most $2d$ from $E(T)$ by deleting from $E(T)$, the appropriate inward and outward edges of $x=x_{1},x_{2},\ldots,x_{d}$, over $\{0,\ldots,n\}$. To output a fixed number of nodes, we assume that $E(T)$ is padded with $2d$ $B$s, $B\gg m$, so that the network can delete the inward and outward edges of $x_{j}\neq 0$ and delete two $B$s corresponding to each $x_{j}=0$. We call such a string a padded Euler string. In this setting, we fix $2n$ output nodes $y=y_{1},\ldots,y_{2n}$ of the network, where $y_{1},\ldots,y_{2n-2d^{\prime}}$, $d^{\prime}\leq d$ is the Euler string of some tree $U$ with the tree edit distance $2d$ from $E(T)$ by deleting $d^{\prime}$ inward and outward edges of non-zero entries of $x$, and the remaining $2d^{\prime}$ entries $y_{2n-2d^{\prime}+1},\ldots,y_{2n}$ are $B$s. An illustration of a TDd-generative ReLU is given in Fig. 7, where $m=5$, $d=3$, $x=1,3,0$, and $T$ is given in Fig. 2 with $E(T)=3,2,7,2,4,9,7,4,9,8$, and the padded $E(T)$ with six $B$s. The network will delete the inward and outward edges of $1$ and $3$ as depicted in the figure, and delete two $B$s corresponding to 0. The resultant string is $y=2,7,4,9,4,9,B,B,B,B$, and by removing $B$s from $y$ we get the desired Euler string $E(U)=2,7,4,9,4,9$ of the tree $U$ as shown in the Fig. 7.
@@
We define TId-generative ReLU as follows. Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)$, and a non-negative integer $d$. We define the TId-generative ReLU to be a ReLU neural network with $4d$ input nodes $x=x_{1},x_{2},\ldots,x_{4d}$ over $\{0,\ldots,n\}$, and $2n+2d$ output nodes $u=u_{1},u_{2},\ldots,u_{2n+2d}$ over $\Sigma$ such that all Euler strings $u$ of trees with the tree edit distance exactly $2d$ from $E(T)$ can be obtained by the insertion of appropriate child nodes $x^{\prime}_{j}$ of $x_{j}$ with inward edges and outward edges of labels $x_{j+3d}$ and $x_{j+3d}+m$, resp., of $x^{\prime}_{j}$, and setting the appropriate ($x_{j+d}$, $x_{j+d}+1$, …, $x_{j+2d}$)-th children of the node $x_{j}$ as the children of $x^{\prime}_{j}$. In this study, we do not consider inserting new nodes as children of a newly inserted node. In this context, $x_{1},\ldots,x_{d}$ represent the nodes for the insertion operations, $x_{1+d},\ldots,x_{d+d}$ and $x_{1+2d},\ldots,x_{d+2d}$ represent the lower and upper bounds to determine the subsequences of children that will be set as the children of the inserted nodes, and $x_{1+3d},\ldots,x_{d+3d}$ represents the labels to be inserted. For convenience, we denote $x_{1},\ldots,x_{d}$ , $x_{1+d},\ldots,x_{2d}$, $x_{1+2d},\ldots,x_{3d}$, and $x_{1+3d},\ldots,x_{4d}$ by $x^{1}$, $x^{2}$, $x^{3}$, and $x^{4}$, respectively. Note that some lower and upper bounds of children may not be valid due to the random nature of $x$, and thus need to be refined to perform appropriate insertion operations. Such invalid bounds, their refinements and appropriate insertions are listed in Table 3, where $D(x_{j}^{1})$ denotes the number of children of the node $x_{j}^{1}$.

### v3

@@ Abstract
The generation of trees with a specified tree edit distance has significant applications across various fields, including computational biology, structured data analysis, and image processing. Recently, generative networks have been increasingly employed to synthesize new data that closely resembles the original datasets. However, the appropriate size and depth of generative networks required to generate data with a specified tree edit distance remain unclear. In this paper, we theoretically establish the existence and construction of generative networks capable of producing trees similar to a given tree with respect to the tree edit distance. Specifically, for a given rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from an alphabet $\Sigma$, and a non-negative integer $d$, we prove that any rooted, ordered, and vertex-labeled tree over $\Sigma$ with tree edit distance at most $d$ from $T$ can be generated using an appropriate random input sequence to a ReLU-based generative network of size $\mathcal{O}(n^{3})$ and constant depth. The proposed networks were implemented and evaluated for generating trees with up to 21 nodes. Due to their deterministic architecture, the networks successfully generated all valid trees within the specified tree edit distance. In contrast, state-of-the-art graph generative models GraphRNN and GraphGDP, which rely on non-deterministic mechanisms, produced significantly fewer valid trees, achieving validation rates of only up to 35% and 48%, respectively. These findings provide a theoretical foundation towards construction of compact generative models and open new directions for exact and valid tree-structured data generation. An implementation of the proposed networks is available at https://github.com/MGANN-KU/TreeGen_ReLUNetworks.
@@ 1 Introduction
As a step towards addressing these issues, we study the exact generation of rooted, ordered, and vertex-labeled trees similar to a given tree using neural networks, and prove the existence of ReLU-based generative networks for the tree edit distance problem. Given a rooted, ordered, and vertex-labeled tree $T$ of size $n+1$ with labels from a symbol set $\Sigma$, we establish the existence of a ReLU-activated network that takes as input a finite random sequence $x$, and outputs a similar tree with edit distance at most $d$ from $T$. The network is deterministic in the sense that, for each random input sequence $x$, it systematically identifies and performs the specified substitution, deletion, and insertion operations to produce a unique output tree $T^{\prime}$ whose tree edit distance from $T$ is at most $d$ (see the proofs of Theorems 1, 2, 3, and 4 for details). Different random input sequences may lead to the same output tree. Nevertheless, by considering an appropriate collection of random input sequences that covers all admissible edit operations, every tree similar to $T$ within distance $d$ can be generated. The key idea of our approach is to first construct a directed, rooted, ordered, and edge-labeled tree based on $T$. We then reduce the tree edit distance problem to a string edit distance problem by representing the tree as an Euler string [48], obtained through a depth-first search (DFS) traversal. The proposed networks are applied on trees with up to 21 nodes to generate similar trees, and are compared with the state-of-the-art graph generative models GraphRNN by You et al. [53] and GraphGDP by Huang et al. [57]. An implementation of the proposed networks is available at https://github.com/MGANN-KU/TreeGen_ReLUNetworks.
The paper is organized as follows: Preliminaries are discussed in Section 2. ReLU generative networks that can identify the indices and labels of the directed edges in the Euler string are discussed in Section 3. Existence of ReLU networks to generate any tree with tree edit distance at most $d$ due to substitution operations is discussed in Section 4. Existence of ReLU networks to generate any tree with edit distance at most (resp., exactly) $d$ due to deletion (resp., insertion) operations is discussed in Section 5. Generation of any tree with tree edit distance at most $d$ due to simultaneous application of deletion, substitution and insertion operations by using a ReLU network is discussed in Section 6. Computational experiments are discussed in Section 7. A conclusion and future directions are given in Section 8. Proofs of some theorems, examples and explanations of the program codes with sample instances are given in Appendix 9.
@@ 2 Preliminaries
Observe that a tree can be completely determined by its Euler string, i.e., $E(T)$ is a canonical representation of $T$ when the labels of roots of the underlying trees is fixed. Therefore any tree edit operations (substitution, deletion, and insertion) on a non-root vertex $u$ of a given tree can be viewed as edit operations on the Euler string of the tree on the entries that correspond to the inward edge $(u,v)$ and the outward edge $(v,u)$ with some refinements to obtain the desired tree. Furthermore, a vertex in $T$ can be specified by its DFS index. Therefore in the rest of the paper, we will use a random sequence $x_{1},x_{2},\ldots,x_{d}$, $d\geq 1$, with integers $x_{j}\in[0,n]$, unless stated otherwise, to specify the DFS indices of the vertices under consideration in a tree with $n+1$ vertices. In this work, we focus on generation of similar trees with the same root as that of the input tree, therefore $x_{j}=0$ is ignored in substitution and deletion operations. Similarly, repeated entries in the case of substitution and deletion operations are ignored. An example random sequence is given in Fig. 2(a). In the rest of the discussion, we call a vertex-labeled or edge-labeled, rooted and ordered tree simply a tree. A list of symbols, variables, and their descriptions used throughout the discussion is provided in Table 3.
@@ 3 Identification of Edge Labels by ReLU Network
We focus on designing generative networks with ReLU as an activation function to generate any tree that are similar to a given tree. More precisely, we are interested in the following problem:
@@
Output: Construct generative networks with ReLU as an activation function that can generate any rooted, ordered, and vertex-labeled tree over $\Sigma$ with tree edit distance at most $d$ from $T$.
@@
Proposition 1 gives a necessary and sufficient condition for the $i$-th entry to be the outward edge of an inward edge at $\ell$-th position in an Euler string $E(T)=t_{1},t_{2},\ldots,t_{2n}$. The condition essentially depends on the number of inward edges and outward edges between the two given positions $i$ and $\ell$. Before going into the details, for any two positions i{i} and $\ell$}, with $1\leq\ell<i\leq 2n$, we denote by $\rm{in}_{\ell,i}$ (resp., $\rm{out}_{\ell,i}$), the total number of inward (resp., outward) edges among the entries $t_{k}$, $\ell<k<i$, where $t_{k}$ is the $k$-th entry of $E(T)$. Consider the tree $T$ given in Fig. 2(a) and its Euler string $E(T)$ given in Fig. 2(c), for $i=7$, $\ell=2$, $\rm{in}_{2,7}=2$ as there are two inward edges $t_{4},t_{5}$ and $\rm{out}_{2,7}=2$ as there are two outward edges $t_{3},t_{6}$ (see Fig. 4(a)). So in this case $\rm{in}_{2,7}=\rm{out}_{2,7}$. Similarly, for $i=9,\ell=3$, $\rm{in}_{3,9}=3>2=\rm{out}_{3,9}$ (see Fig. 4(b)), for $i=9,\ell=5$, $\rm{in}_{5,9}=1<2=\rm{out}_{5,9}$ (see Fig. 4(c)), and for $i=7,\ell=4$, $\rm{in}_{4,7}=1=\rm{out}_{4,7}$ (see Fig. 4(d)).
Figure 4: (a)-(c)An illustration of the number of inward edges (green) (resp., outward edges (blue)) for the edges $t_{i},i=7,9$ and $t_{\ell},\ell=2,3,5$ which are depicted in gray boxes; (d) An example in which Proposition 1 holds.
@@
Consider the tree $T$ given in Fig. 2(a) and its Euler string $E(T)$ shown in Fig. 2(c). We want to determine if the edges $t_{i}$, $i=4,7,9$ in $E(T)$ are outward edges of some inward edges by using Proposition 1. We see that Proposition 1(ii) does not hold for $i=4$ and any $\ell\in[1,3]$, therefore $t_{4}$ is not an outward edge of any inward edge, which is consistent with the fact. For $i=7$, Proposition 1(ii) and (iii) are satisfied for both $\ell=2,4$. For $\ell=2$ (resp., $\ell=4$), $\rm{in}_{2,7}=2=\rm{out}_{2,7}$ (resp., $\rm{in}_{4,7}=1=\rm{out}_{4,7}$) (see Fig. 4(a)). That is $t_{2}$ and $t_{4}$ are both candidate inward edges of $t_{7}$. By using Proposition 1(iv), $t_{7}$ is the outward edge of $t_{4}$(see Fig. 4(d)). For $i=9$, we see that $t_{5}$ and $t_{8}$ are the only edges that satisfy Proposition 1(i)-(ii), but $t_{5}$ does not satisfy the Proposition 1(iii). Therefore $t_{8}$ is the inward edge of $t_{9}$ implying that $t_{9}$ is an outward edge.
@@ 4 TSd-generative ReLU
Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)=t_{1},t_{2},\ldots,t_{2n}$, and a non-negative integer $d$. We define the TSd-generative ReLU to be a ReLU neural network with $2d$ input nodes $x=x_{1},x_{2},\ldots,x_{2d}$ over $\{0,\ldots,n\}$, and $2n$ output nodes $u=u_{1},u_{2},\ldots,u_{2n}$ over $\Sigma$ such that all Euler strings $u$ of trees with the tree edit distance at most $2d$ from $E(T)$ can be obtained by the substitution of appropriate $x_{j+d}$ and $x_{j+d}+m$ at the inward edge and outward edge of $x_{j}\neq 0$, respectively. In this context $x_{1},\ldots,x_{d}$ represents the inward edges for the substitution operations, while $x_{d+1},\ldots,x_{2d}$ denotes the values to be substituted during these operations. The generated Euler strings have distance at most $2d$ as substitution at root and repeated terms in $x$ will be ignored. An illustration of such a network is given in Fig. 6, where $m=5$, $d=3$, $x=1,3,0,5,1,2$, means that the labels of the inward edges (resp., outward edges) of $x_{1}=1,x_{2}=3$ are substituted by $x_{4}=5,x_{5}=1$ (resp., $10,6$). Since the entry $x_{3}=0$ and $x_{6}=2$ corresponds to the root and its label, resp., and so they are ignored.
@@ 5 TDd, TId-generative ReLU
Let $T$ be a tree with $n+1$ nodes, labels from $\Sigma=\{1,2,\ldots,m\}$ and Euler string $E(T)=t_{1},t_{2},\ldots,t_{2n}$, and a non-negative integer $d$. We define the TDd-generative ReLU to be a ReLU neural network that can generate all Euler strings over $\Sigma$ with the tree edit distance at most $2d$ from $E(T)$ by deleting from $E(T)$, the appropriate inward and outward edges of $x=x_{1},x_{2},\ldots,x_{d}$, over $\{0,\ldots,n\}$. The generated Euler strings have distance at most $2d$ as deletion of root and repeated terms in $x$ will be ignored. Since the network does not always delete exactly $2d$ elements from the input Euler string, while a neural network must have a fixed number of nodes in each layer, we pad $E(T)$ with $2d$ number of $B$s where $B\gg m$, to ensure a fixed output dimension. To delete a vertex, both its inward and outward edges must be deleted. Therefore, for each $x_{j}\neq 0$, the network deletes two elements of the input Euler string corresponding to the inward and outward edges of $x_{j}$, and for each $x_{j}=0$, it deletes two padding symbols $B$. We call such a string a padded Euler string We call such a string a padded Euler string. In this setting, we fix $2n$ output nodes $y=y_{1},\ldots,y_{2n}$ of the network, where $y_{1},\ldots,y_{2n-2d^{\prime}}$, $d^{\prime}\leq d$ is the Euler string of some tree $U$ with the tree edit distance $2d$ from $E(T)$ by deleting $d^{\prime}$ inward and outward edges of non-zero entries of $x$, and the remaining $2d^{\prime}$ entries $y_{2n-2d^{\prime}+1},\ldots,y_{2n}$ are $B$s. An illustration of a TDd-generative ReLU is given in Fig. 7, where $m=5$, $d=3$, $x=1,3,0$, and $T$ is given in Fig. 2 with $E(T)=3,2,7,2,4,9,7,4,9,8$, and the padded $E(T)$ with six $B$s. The network will delete the inward and outward edges of $1$ and $3$ as depicted in the figure, and delete two $B$s corresponding to 0. The resultant string is $y=2,7,4,9,4,9,B,B,B,B$, and by removing $B$s from $y$ we get the desired Euler string $E(U)=2,7,4,9,4,9$ of the tree $U$ as shown in the Fig. 7.
@@
We define TId-generative ReLU as follows. Let $T$ be a tree with $n+1$ nodes and labels from $\Sigma=\{1,2,\ldots,m\}$, Euler string $E(T)$, and a non-negative integer $d$. We define the TId-generative ReLU to be a ReLU neural network with $4d$ input nodes $x=x_{1},x_{2},\ldots,x_{4d}$ over $\{0,\ldots,n\}$, and $2n+2d$ output nodes $u=u_{1},u_{2},\ldots,u_{2n+2d}$ over $\Sigma$ such that all Euler strings $u$ of trees with the tree edit distance exactly $2d$ from $E(T)$ can be obtained by the insertion of appropriate child nodes $x^{\prime}_{j}$ of $x_{j}$ with inward edges and outward edges of labels $x_{j+3d}$ and $x_{j+3d}+m$, resp., of $x^{\prime}_{j}$, and setting the appropriate ($x_{j+d}$, $x_{j+d}+1$, $x_{j+d}+2$, …, $x_{j+2d}$)-th children of the node $x_{j}$ as the children of $x^{\prime}_{j}$. It means $x^{\prime}_{j}$ becomes the parent of a sequence of children that starts from $x_{j+d}$ and ends at $x_{j+2d}$. In this study, we do not consider inserting new nodes as children of a newly inserted node. In this context, $x_{1},\ldots,x_{d}$ represent the nodes for the insertion operations, $x_{1+d},\ldots,x_{d+d}$ and $x_{1+2d},\ldots,x_{d+2d}$ represent the lower and upper bounds to determine the subsequences of children that will be set as the children of the inserted nodes, and $x_{1+3d},\ldots,x_{d+3d}$ represents the labels to be inserted. For convenience, we denote $x_{1},\ldots,x_{d}$ , $x_{1+d},\ldots,x_{2d}$, $x_{1+2d},\ldots,x_{3d}$, and $x_{1+3d},\ldots,x_{4d}$ by $x^{1}$, $x^{2}$, $x^{3}$, and $x^{4}$, respectively. Note that lower and upper bounds on the number of children must not exceed the total number of children of a node $x_{j}^{1}$. Let $D(x_{j}^{1})$ denote the total number of children of a node $x_{j}^{1}$ then we require $x_{j}^{2},x_{j}^{3}\leq D(x_{j}^{1})$. Due to the random nature of $x$, some of the lower and upper bounds of children may not be valid, and thus need to be refined to perform appropriate insertion operations. Such invalid bounds, their refinements and appropriate insertions are listed in Table 3, where $D(x_{j}^{1})$ denotes the number of children of the node $x_{j}^{1}$.